# The Role of the Schwinger Effect in Superradiant Axion Lasers

### v2

@@ III.2 Positron/Electron Escape Rate
[…]
@@
$\Delta r-\rho_{0}$ $=T_{e^{\pm}}\sqrt{1-\frac{16m_{e}^{2}V^{2}}{9\sigma_{T}^{2}m_{\phi}^{2}N_{\gamma}^{2}T_{e^{\pm}}^{2}}}-\frac{4m_{e}V}{3\sigma_{T}m_{\phi}N_{\gamma}}\tan^{-1}\left(\sqrt{\frac{9\sigma_{T}^{2}m_{\phi}^{2}N_{\gamma}^{2}T_{e^{\pm}}^{2}}{16m_{e}^{2}V^{2}}-1}\right)$ (30)
@@
as a dimensionless parameter for the behavior of $\Gamma_{e^{\pm}}$. $\Gamma_{e^{\pm}}$ increases monotonically as $\chi$ increases, asymptotically approaching $\Gamma_{\gamma}$. This behavior is analyzed in greater detail in the appendix, and it is plotted in Fig. 1a.
Figure 1: Plot of functions of $\chi$, which parameterizes how relativistic the electrons and positrons become as a result of radiation pressure. (a) demonstrates the behavior of $\Gamma_{e^{\pm}}$. At low $\chi$, there is negligible radiation pressure driving electrons and positrons to exit the $2p$-cloud, and so $\Gamma_{e^{\pm}}\to 0$. For large $\chi$ the electrons and positrons are quickly accelerated to very close to the speed of light, causing them to exit the $2p$-cloud at roughly the same rate as the photons, and as a result $\Gamma_{e^{\pm}}\to\Gamma_{\gamma}$. (b) demonstrates the behavior of $\Gamma_{\textrm{ann}}$. Slower particles present a larger cross-section, and therefore pair annihilation occurs frequently for small $\chi$ but vanishes as $\chi\to\infty$. Note that $\Gamma_{\textrm{ann}}$ is the rate of annihilation per number of pairs squared, not the total rate; in practice, large $\chi$ will correlate with large $N_{e^{+}e^{-}}$, so the total rate of annihilation will increase with $\chi$. (c) demonstrates the behavior of $g(\chi)$, which scales the number of electron-positron pairs at which annihilation becomes relevant. We show in the appendix that, for large $\chi$, $g(\chi)$ grows as $\frac{\chi^{2}}{(\ln\chi)^{2}}$.
@@ III.3 Pair Annihilation
[…]
@@
We show in the appendix that $\Gamma_{\textrm{ann}}V$ is a function only of $\chi$, which is displayed in Fig. 1b. As $\chi$ increases, the electrons and positrons become more relativistic, and so both the cross section and the rate of annihilation decrease.
@@ IV Boltzmann Equations and Parameter Space
Initially, the rate of pair production is vanishingly small, so $N_{\phi}$ and $N_{\gamma}$ behave identically to the description given in [23]. When $N_{\gamma}$ gets within a few orders of magnitude of $N_{\gamma}^{\textrm{Schw}}$ (roughly $10^{-4}N_{\gamma}^{\textrm{Schw}}$), electron-positron pairs begin to be produced, and these electron-positron pairs impact the axion decay rate. Thus we need only consider the case when $N_{\gamma}$ is approaching $N_{\gamma}^{\textrm{Schw}}$.
@@
$N_{e^{+}e^{-}}^{\textrm{Schw}}=8.8*10^{48}\left(\frac{\alpha}{0.037}\right)^{-4}\left(\frac{m_{\phi}}{10^{-5}\textrm{ eV}}\right)^{-4}.$ (48)
@@
which uses the hat notation defined in the appendix. $g(\chi)$ is plotted in Fig. 1c, and we show in the appendix that its asymptotic form is
@@
where $\Gamma_{\textrm{Schw}}^{\textrm{eq}}$ and $\Gamma_{e^{\pm}}^{\textrm{eq}}$ represent $\Gamma_{\textrm{Schw}}$ and $\Gamma_{e^{\pm}}$ evaluated at $N_{\gamma}^{\textrm{eq}}$, and $N_{\gamma}^{\textrm{eq,w/o}}=\frac{\Gamma_{s}}{C_{1}\Gamma_{\phi}}$ is the expression for $N_{\gamma}^{\textrm{eq}}$ calculated without accounting for the Schwinger effect. This equation does not have a solution for all regions of parameter space; for some values of $m_{\phi}$ and $\alpha$, it is simply impossible for the system to reach equilibrium. This represents the case where the suppression of axion decay is so great that the photons will never build up to the point where they come into equilibrium with the axions. While the number of photons will reach or approach $N_{\gamma}^{\textrm{ceiling}}$, the number of axions will continue growing unbounded. This is reflected mathematically by the fact that $C_{1}\Gamma^{\prime}_{\phi}N_{\gamma}$ has some maximum attainable value, and, if this maximum value is less than $\Gamma_{s}$, then $N_{\phi}$ will increase without bound. This will eventually result in the black hole spinning down to the point where superradiance ceases. These regions of parameter space where the system is unstable are shown on Fig. 2.
It should be noted that, while this paper has focused on axion decay, which cannot happen when the plasma frequency exceeds $\frac{m_{\phi}}{2}$, it has been shown that axion clouds may still emit light after exceeding this limit, via axion annihilation, i.e. $\phi+\phi\to\gamma+\gamma$ [24, 27]. Indeed, it is possible to have any number of photons annihilate to produce two photons, $n\phi\to\gamma+\gamma$, with each individual annihilation process shutting off when $\omega_{p}>n\frac{m_{\phi}}{2}$. These higher-order processes will become more relevant in the unstable case, as the axion cloud is expected to reach greater densities than is possible outside of the unstable case, and these processes may in fact allow the system to reach equilibrium (making the term “unstable case” something of a misnomer; it is only unstable in this particular analysis, which ignores higher-level processes). This, however, is beyond the scope of this paper.
[…]
@@
Figure 3: Plot of the enhancement of the equilibrium luminosity against $\alpha$ and $m_{\phi}$. $L^{\textrm{eq}}$ is the equilibrium luminosity, while $L^{\textrm{eq,w/o}}$ is the equilibrium luminosity found in [23], i.e. ignoring the effects of Schwinger pairs. The same unstable region shown in Fig. 2 is also visible here. In the part of parameter space where an equilibrium state exists, we can see that the equilibrium luminosity is hardly enhanced at all, which means that the role of Schwinger production is negligible to the final state of the BLAST. However, close to the unstable region, there is a region where the enhancement becomes more significant. Note that computational constraints limit the resolution of the section of the plot with the greatest enhancement. This means that, even though the largest enhancement displayed on this plot is .02.02.02.02, it is likely that even greater enhancements can be achieved.
@@ V Conclusion
An important caveat to these results is that our calculation of the rate of pair production involved two significant approximations: we ignored the role of both the magnetic field and the axion field. Therefore, all numerical results herein must be viewed as approximations, accurate to within a few orders of magnitude. A more careful analysis would require the inclusion of both the magnetic field and the axion’s enhancement to the Schwinger effect, as described in [47]. Additionally, we have assumed that all particles are uniformly distributed throughout the $2p$-cloud; however, the simulations conducted in [27] have demonstrated that this approximation can fail to predict significant phenomena. It is therefore necessary to do away with this assumption as well, which will entail replacing the analytical calculations in this paper with numerical simulations.
@@ Acknowledgements
The author thanks Devin Walker and João Rosa for their helpful feedback. This work was supported by the National Science Foundation under grant OIA-2033382.
Appendix A Behavior of functions of $\chi$

### v3

@@ III.2 Positron/Electron Escape Rate
and therefore for all superradiant modes ($\alpha\lesssim.5$), $r\gg r_{s}$ at all points within the $2p$-cloud. We may therefore neglect this prefactor; put another way, the $2p$-cloud is far enough away from the black hole that we may treat spacetime as flat.
@@
$\Delta r-\rho_{0}$ $=\sqrt{T_{e^{\pm}}\left(\frac{8Vm_{e}}{3\sigma_{T}m_{\phi}N_{\gamma}}+T_{e^{\pm}}\right)}$
$\hskip 28.45274pt+\frac{4Vm_{e}}{3\sigma_{T}m_{\phi}N_{\gamma}}\left(\frac{\pi}{3}-2\tan^{-1}\frac{\sqrt{3}\left(\frac{4Vm_{e}}{3\sigma_{T}m_{\phi}N_{\gamma}}+T_{e^{\pm}}\right)+2\sqrt{T_{e^{\pm}}\left(\frac{8Vm_{e}}{3\sigma_{T}m_{\phi}N_{\gamma}}+T_{e^{\pm}}\right)}}{\frac{4Vm_{e}}{\sigma_{T}m_{\phi}N_{\gamma}}+T_{e^{\pm}}}\right).$ (38)
@@
as a dimensionless parameter for the behavior of $\Gamma_{e^{\pm}}$. $\Gamma_{e^{\pm}}$ increases monotonically as $\chi$ increases, asymptotically approaching $\Gamma_{\gamma}$. This behavior is analyzed in greater detail in Appendix B, and it is plotted in Fig. 1a.
Figure 1: Plot of functions of $\chi$, which parameterizes how relativistic the electrons and positrons become as a result of radiation pressure. (a) demonstrates the behavior of $\Gamma_{e^{\pm}}$. At low $\chi$, there is negligible radiation pressure driving electrons and positrons to exit the $2p$-cloud, and so $\Gamma_{e^{\pm}}\to 0$. For large $\chi$ the electrons and positrons are quickly accelerated to very close to the speed of light, causing them to exit the $2p$-cloud at roughly the same rate as the photons, and as a result $\Gamma_{e^{\pm}}\to\Gamma_{\gamma}$. (b) demonstrates the behavior of $\Gamma_{\textrm{ann}}$. Slower particles present a larger cross-section, and therefore pair annihilation occurs frequently for small $\chi$ but vanishes as $\chi\to\infty$. Note that $\Gamma_{\textrm{ann}}$ is the rate of annihilation per number of pairs squared, not the total rate; in practice, large $\chi$ will correlate with large $N_{e^{+}e^{-}}$, so the total rate of annihilation will increase with $\chi$. (c) demonstrates the behavior of $g(\chi)$, which scales the number of electron-positron pairs at which annihilation becomes relevant. We show in Appendix B that, for large $\chi$, $g(\chi)$ grows as $\frac{\chi^{2}}{(\ln\chi)^{2}}$.
@@ III.3 Pair Annihilation
$\frac{\textrm{d}n_{e^{\pm}}}{\textrm{d}t}=-n_{e^{+}}n_{e^{-}}\int dv_{e^{+}}$ $dv_{e^{-}}f(v_{e^{+}})f(v_{e^{-}})$
$\times\lvert v_{e^{+}}-v_{e^{-}}\rvert\sigma(v_{\textrm{com}}).$ (44)
@@
We show in Appendix B that $\Gamma_{\textrm{ann}}V$ is a function only of $\chi$, which is displayed in Fig. 1b. As $\chi$ increases, the electrons and positrons become more relativistic, and so both the cross section and the rate of annihilation decrease.
@@ IV Boltzmann Equations and Parameter Space
Note that, because we are modeling the total number of particles in the axion cloud, we have only included phenomena that can introduce or remove particles from the cloud. There are likely to be additional phenomena that alter the distribution of particles within the axion cloud (e.g. scattering between electrons and positrons), but these are not relevant to this paper.
Numerical solutions to these equations, for two different sets of parameters, are shown in Fig. 2. Initially, the rate of pair production is vanishingly small, so $N_{\phi}$ and $N_{\gamma}$ behave identically to the description given in [23]. When $N_{\gamma}$ gets within a few orders of magnitude of $N_{\gamma}^{\textrm{Schw}}$ (roughly $10^{-4}N_{\gamma}^{\textrm{Schw}}$), electron-positron pairs begin to be produced, and these electron-positron pairs impact the axion decay rate. Thus we need only consider the case when $N_{\gamma}$ is approaching $N_{\gamma}^{\textrm{Schw}}$.
    (ai)                        (aii)
    (bi)                        (bii)
Figure 2: Numerical simulations of the growth of $N_{\phi}$, $N_{\gamma}$, and $N_{e^{+}e^{-}}$. (a) is done with $\alpha=.037$ while (b) is done with $\alpha=.02$, with the other parameters set to $m_{\phi}=10^{-5}\textrm{ eV}$, $\tilde{a}=.7$, and $K=1$. (aii) and (bii) show zoomed-in views of the onset of lasing, when $N_{\gamma}$ increases rapidly. In both cases, $N_{e^{+}e^{-}}$ increases alongside $N_{\gamma}$, matching with the approximation $N_{e^{+}e^{-}}\approx\frac{\Gamma_{\textrm{Schw}}}{\Gamma_{e^{\pm}}}$. Prior to the onset of lasing, $N_{\phi}$ and $N_{\gamma}$ increase exponentially, while $N_{e^{+}e^{-}}$ is negligible. In (a), $N_{e^{+}e^{-}}$ approaches $N_{e^{+}e^{-}}^{\textrm{shutoff}}$, slowing axion decay, so that $N_{\gamma}$ stops increasing while $N_{\phi}$ continues increasing. In (b), $N_{e^{+}e^{-}}$ is not large enough to significantly affect axion decay, and so $N_{\gamma}$ and $N_{\phi}$ settle into their equilibrium values. (a) lies in the unstable region of parameter space, while (b) lies in the unenhanced region (see Figs. 3 and 4). Note that, in both plots, $N_{\gamma}$ always remains multiple orders of magnitude below $N_{\gamma}^{\textrm{Schw}}$, allowing for the approximation given in eq. 28.
@@
$N_{e^{+}e^{-}}^{\textrm{Schw}}=4*10^{48}\left(\frac{\alpha}{0.037}\right)^{-4}\left(\frac{m_{\phi}}{10^{-5}\textrm{ eV}}\right)^{-4}.$ (56)
@@
which uses the hat notation defined in Appendix B. $g(\chi)$ is plotted in Fig. 1c, and we show in Appendix B that its asymptotic form is
@@
Figure 3: Plot of the $\alpha m_{\phi}$ plane of parameter space. The other parameters were taken to be $\tilde{a}=.7$ and $K=1$. For larger values of $\alpha$ and smaller values of $m_{\phi}$, the system is unstable, represented by the black-colored region of parameter space. In this region, the electron-positron plasma suppresses the axion decay so much that the photon number never becomes large enough to halt the growth of the axion number. The axion number therefore continues to grow superradiantly, causing the black hole to spin down until it is no longer superradiant. In the colored regions of parameter space, it is possible for the system to reach equilibrium, and the equilibrium luminosity is plotted. Especially for light axions, these luminosities are extremely high, making BLASTs potentially significant for future experiments. In the shaded region of parameter space, $L^{\textrm{eq}}$ depends on both $\alpha$ and $m_{\phi}$, but the dependence on $\alpha$ is too small to see on this plot.
Figure 4: Plot of the enhancement of the equilibrium luminosity against $\alpha$ and $m_{\phi}$. $L^{\textrm{eq}}$ is the equilibrium luminosity, while $L^{\textrm{eq,w/o}}$ is the equilibrium luminosity found in [23], i.e. ignoring the effects of Schwinger pairs. The same unstable region shown in Fig. 2 is also visible here. In the part of parameter space where an equilibrium state exists, we can see that the equilibrium luminosity is hardly enhanced at all, which means that the role of Schwinger production is negligible to the final state of the BLAST. However, close to the unstable region, there is a region where the enhancement becomes more significant. Note that computational constraints limit the resolution of the section of the plot with the greatest enhancement. This means that, even though the largest enhancement displayed on this plot is .02.02, it is likely that even greater enhancements can be achieved.
@@
where $\Gamma_{\textrm{Schw}}^{\textrm{eq}}$ and $\Gamma_{e^{\pm}}^{\textrm{eq}}$ represent $\Gamma_{\textrm{Schw}}$ and $\Gamma_{e^{\pm}}$ evaluated at $N_{\gamma}^{\textrm{eq}}$, and $N_{\gamma}^{\textrm{eq,w/o}}=\frac{\Gamma_{s}}{C_{1}\Gamma_{\phi}}$ is the expression for $N_{\gamma}^{\textrm{eq}}$ calculated without accounting for the Schwinger effect. This equation does not have a solution for all regions of parameter space; for some values of $m_{\phi}$ and $\alpha$, it is simply impossible for the system to reach equilibrium. This represents the case where the suppression of axion decay is so great that the photons will never build up to the point where they come into equilibrium with the axions. While the number of photons will reach or approach $N_{\gamma}^{\textrm{ceiling}}$, the number of axions will continue growing unbounded. This is reflected mathematically by the fact that $C_{1}\Gamma^{\prime}_{\phi}N_{\gamma}$ has some maximum attainable value, and, if this maximum value is less than $\Gamma_{s}$, then $N_{\phi}$ will increase without bound. This will eventually result in the black hole spinning down to the point where superradiance ceases. These regions of parameter space where the system is unstable are shown in Fig. 3.
It should be noted that, while this paper has focused on axion decay, which cannot happen when the plasma frequency exceeds $\frac{m_{\phi}}{2}$, it has been shown that axion clouds may still emit light after exceeding this limit, via axion annihilation, i.e. $\phi+\phi\to\gamma+\gamma$ [24, 27]. Indeed, it is possible to have any number of axions annihilate to produce two photons, $n\phi\to\gamma+\gamma$, with each individual annihilation process shutting off when $\omega_{p}>n\frac{m_{\phi}}{2}$. These higher-order processes will become more relevant in the unstable case, as the axion cloud is expected to reach greater densities than is possible outside of the unstable case, and these processes may in fact allow the system to reach equilibrium (making the term “unstable case” something of a misnomer; it is only unstable in this particular analysis, which ignores higher-level processes). This, however, is beyond the scope of this paper.
@@ V Conclusion
An important caveat to these results is that our calculation of the rate of pair production ignored the phenomenon of axion assistance to the Schwinger effect [44, 45]. Because a BLAST occurs in an axion field, we should expect the rate of pair production to be greater than what is calculated using the standard Schwinger effect (i.e., based only on the electromagnetic field, without considering the axion field). The rate of Schwinger production used in this paper therefore serves as a lower bound. The closeness this lower bound to reality will depend on the axion-electron coupling, $g_{ae}$, with the true rate of pair production approaching the rate used in this paper only in the limit $g_{ae}\to 0$. We may predict that increasing $g_{ae}$ will have the effect of expanding the unbounded region of parameter space, as it becomes easier for the Schwinger effect to create a plasma of sufficient density that axion decay is shut off. However, future research is needed to examine, in greater detail, the behavior of BLASTs in the case of $g_{ae}>0$.
It is also worth noting that we have assumed that all particles are uniformly distributed throughout the $2p$-cloud; however, the simulations conducted in [27] have demonstrated that this approximation can fail to predict significant phenomena. It is therefore necessary to do away with this assumption as well, which will entail replacing the analytical calculations in this paper with numerical simulations.
@@ Acknowledgements
The author thanks Devin Walker and João Rosa for their valuable feedback, as well as Roberto Onofrio and Jens Mahlmann for helpful conversations. This work was supported by the National Science Foundation under grant OIA-2033382.
The data that support the findings of this article are openly available [56].
Appendix A Calculation of the rate of pair production
In this appendix we perform the integral in eq. 24 in order to attain an expression for $\Gamma_{\textrm{Schw}}$. We begin by rewriting our integral in terms only of $E_{\parallel}$, $B_{\parallel}$, and $\theta$:
$\Gamma_{\textrm{Schw}}=\frac{e^{2}V}{4\pi^{3}\sigma_{\textrm{EM}}^{6}}\int_{0}^{\pi}\int_{0}^{\infty}\int_{0}^{\infty}\textrm{d}\theta$ $\textrm{d}E_{\parallel}\textrm{d}B_{\parallel}E_{\parallel}^{3}B_{\parallel}^{3}\left(E_{\parallel}^{2}+B_{\parallel}^{2}\right)$
$\frac{\sin\theta\lvert\sec^{3}\theta\rvert e^{-\frac{\sqrt{\left(E_{\parallel}^{2}-B_{\parallel}^{2}\right)^{2}+4E_{\parallel}^{2}B_{\parallel}^{2}\sec^{2}\theta}}{2\sigma_{\textrm{EM}}^{2}}}}{\sqrt{\left(E_{\parallel}^{2}-B_{\parallel}^{2}\right)^{2}+4E_{\parallel}^{2}B_{\parallel}^{2}\sec^{2}\theta}}\sum_{n=1}^{\infty}\frac{1}{n}\coth\frac{n\pi B_{\parallel}}{E_{\parallel}}e^{-n\pi\frac{E_{c}}{E_{\parallel}}}.$ (61)
The integral over $\theta$ may be performed through the substitution $u=\sqrt{\left(E_{\parallel}^{2}-B_{\parallel}^{2}\right)^{2}+4E_{\parallel}^{2}B_{\parallel}^{2}\sec^{2}\theta}$, yielding
$\Gamma_{\textrm{Schw}}=\frac{e^{2}V}{4\pi^{3}\sigma_{\textrm{EM}}^{4}}\int_{0}^{\infty}\int_{0}^{\infty}\textrm{d}E_{\parallel}\textrm{d}B_{\parallel}E_{\parallel}B_{\parallel}\left(E_{\parallel}^{2}+B_{\parallel}^{2}\right)e^{-\frac{E_{\parallel}^{2}+B_{\parallel}^{2}}{2\sigma_{\textrm{EM}}^{2}}}\sum_{n=1}^{\infty}\frac{1}{n}\coth\frac{n\pi B_{\parallel}}{E_{\parallel}}e^{-n\pi\frac{E_{c}}{E_{\parallel}}}.$ (62)
We may expand $\coth$ as a series of exponentials, allowing us to rewrite this as
$\Gamma_{\textrm{Schw}}=\frac{e^{2}V}{4\pi^{3}\sigma_{\textrm{EM}}^{4}}\sum_{n=1}^{\infty}\frac{1}{n}\int_{0}^{\infty}\textrm{d}E_{\parallel}E_{\parallel}e^{-\frac{E_{\parallel}^{2}}{2\sigma_{\textrm{EM}}^{2}}}e^{-n\pi\frac{E_{c}}{E_{\parallel}}}\sum_{k=0}^{\infty}\left(2-\delta_{k0}\right)\int_{0}^{\infty}\textrm{d}B_{\parallel}B_{\parallel}\left(E_{\parallel}^{2}+B_{\parallel}^{2}\right)e^{-\frac{B_{\parallel}^{2}}{2\sigma_{\textrm{EM}}^{2}}}e^{-2nk\pi\frac{B_{\parallel}}{E_{\parallel}}}.$ (63)
[…]
This identity requires a number of conditions for its validity, but the only ones relevant to this discussion are that $\sigma$ and $\omega$ must both be nonzero; consequently, we must separate the $k=0$ term from the $k\neq 0$ terms, applying the above identity only to the latter. This yields
$\Gamma_{\textrm{Schw}}$ $=\frac{e^{2}V}{4\pi^{3}\sigma_{\textrm{EM}}^{4}}\sum_{n=1}^{\infty}\frac{1}{n}\int_{0}^{\infty}\textrm{d}E_{\parallel}E_{\parallel}e^{-\frac{E_{\parallel}^{2}}{2\sigma_{\textrm{EM}}^{2}}}e^{-n\pi\frac{E_{c}}{E_{\parallel}}}\Bigg(2\sigma_{\textrm{EM}}^{4}+\sigma_{\textrm{EM}}^{2}E_{\parallel}^{2}$
[…]
Defining $v=E_{\parallel}^{2}$:
$\Gamma_{\textrm{Schw}}$ $=\frac{e^{2}V}{8\pi^{3.5}\sigma_{\textrm{EM}}^{4}}\sum_{n=1}^{\infty}\frac{1}{n}\int_{0}^{\infty}\textrm{d}ve^{-\frac{v}{2\sigma_{\textrm{EM}}^{2}}}G^{0,2}_{2,0}\left(\frac{4v}{n^{2}\pi^{2}E_{c}^{2}}\left\lvert\begin{tabular}[]{cc}$1,\frac{1}{2}$\\ $-$\end{tabular}\right.\right)\Bigg(2\sigma_{\textrm{EM}}^{4}+\sigma_{\textrm{EM}}^{2}v$ (76)
[…]
The first two terms may be integrated using eq. 68. Integrating the higher-order terms requires the use of the bivariate Meijer G-function [58, 59, 60], defined as
$G^{m_{1},n_{1}:m_{2},n_{2};m_{3},n_{3}}_{\,p_{1},\,q_{1}\,:\,p_{2},\,q_{2}\,;\,p_{3},\,q_{3}}\left(x,y\left\lvert\begin{tabular}[]{ccccc}$\left(\alpha^{(1)}_{p_{1}}\right)$&:&$\left(\alpha^{(2)}_{p_{2}}\right)$&;&$\left(\alpha^{(3)}_{p_{3}}\right)$\\ $\left(\beta^{(1)}_{q_{1}}\right)$&:&$\left(\beta^{(2)}_{q_{2}}\right)$&;&$\left(\beta^{(3)}_{q_{3}}\right)$\end{tabular}\right.\right)=\frac{-1}{4\pi^{2}}\int_{L_{1}}\int_{L_{2}}\Psi_{1}(s+t)\Psi_{2}(s)\Psi_{3}(t)x^{s}y^{t}\textrm{d}s\textrm{d}t,$ (84)
where $L_{1}$ and $L_{2}$ are suitable contours and
$\Psi_{i}(u)=\frac{\displaystyle\prod_{j=1}^{m_{i}}\Gamma\left(\beta^{(i)}_{j}-u\right)\prod_{j=1}^{n_{i}}\Gamma\left(1-\alpha^{(i)}_{j}+u\right)}{\displaystyle\prod_{j=m_{i}+1}^{q_{i}}\Gamma\left(1-\beta^{(i)}_{j}+u\right)\prod_{j=n_{i}+1}^{p_{i}}\Gamma\left(\alpha^{(i)}_{j}-u\right)}.$ (85)
When $p_{1}=q_{1}=0$, the bivariate Meijer G-function becomes a product of two univariate Meijer G-functions. The integral may then be performed using the identity [61]
[…]
yielding
[…]
With some manipulation, we may rewrite this as
[…]
We now wish to evaluate this in the limit $\sigma_{\textrm{EM}}\ll E_{c}$. To do this, we first rewrite the bivariate Meijer G-function as
[…]
$\hskip 85.35826pt=\frac{1}{2\pi i}\int_{L}\Gamma(t)\Gamma\left(\frac{1}{2}+t\right)\Gamma(1-t)G^{0,4}_{4,1}\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\left\lvert\begin{tabular}[]{c}$1,\frac{1}{2},-2,-t$\\ $-1$\end{tabular}\right.\right)\left(\frac{1}{n^{2}k^{2}\pi^{2}}\right)^{t}\textrm{d}t.$ (114)
In the aforementioned limit,
$G^{0,4}_{4,1}\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\left\lvert\begin{tabular}[]{c}$1,\frac{1}{2},-2,-t$\\ $-1$\end{tabular}\right.\right)\approx\frac{2\pi}{\sqrt{3}}e^{-3\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{\frac{-1}{3}}}\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{-\frac{1}{2}-\frac{t}{3}},$ (115)
and thus we get
[…]
$\hskip 85.35826pt\approx\frac{n\pi^{2}E_{c}}{\sqrt{6}\sigma_{\textrm{EM}}}e^{-3\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{\frac{-1}{3}}}\frac{1}{2\pi i}\int_{L}\Gamma(t)\Gamma\left(\frac{1}{2}+t\right)\Gamma(1-t)\left(\frac{1}{k^{2}}\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}\right)^{t}\textrm{d}t.$ (119)
The sum over $k$ becomes a Riemann zeta function:
[…]
$\hskip 28.45274pt=\frac{n\pi^{2}E_{c}}{\sqrt{6}\sigma_{\textrm{EM}}}e^{-3\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{\frac{-1}{3}}}\frac{1}{2\pi i}\int_{L}\Gamma(t)\Gamma\left(\frac{1}{2}+t\right)\Gamma(1-t)\left(\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}\right)^{t}\zeta(2t)\textrm{d}t$
$\hskip 28.45274pt=\frac{n\pi^{2}E_{c}}{\sqrt{6}\sigma_{\textrm{EM}}}e^{-3\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{\frac{-1}{3}}}\frac{1}{2\pi i}\int_{0}^{\infty}\frac{x^{-1}\textrm{d}x}{e^{x}-1}\int_{L}\frac{\Gamma(t)\Gamma\left(\frac{1}{2}+t\right)}{\Gamma(2t)}\Gamma(1-t)\left(x^{2}\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}\right)^{t}\textrm{d}t,$ (123)
where we have used the zeta function’s integral representation,
$\zeta(t)=\sum_{k=1}^{\infty}\frac{1}{k^{t}}=\frac{1}{\Gamma(t)}\int_{0}^{\infty}\frac{x^{t-1}}{e^{x}-1}\textrm{d}x.$ (124)
With some manipulation, the inner integral may be rewritten as a Meijer G-function with a known form, yielding
[…]
$\hskip 28.45274pt=\frac{n\pi^{2}E_{c}}{\sqrt{6}\sigma_{\textrm{EM}}}e^{-3\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{\frac{-1}{3}}}\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}\frac{\sqrt{\pi}}{2}\int_{0}^{\infty}\frac{x\textrm{d}x}{e^{x}-1}e^{-\frac{x^{2}}{4}\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}}.$ (128)
Our expression for the total rate of pair production then becomes
$\Gamma_{\textrm{Schw}}=\frac{e^{2}E_{c}V\sigma_{\textrm{EM}}}{2\sqrt{6}\pi^{1.5}}\sum_{n=1}^{\infty}e^{-3\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{\frac{-1}{3}}}\left(1+\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}\int_{0}^{\infty}\frac{x\textrm{d}x}{e^{x}-1}e^{-\frac{x^{2}}{4}\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}}\right).$ (129)
Terms in this summation vanish as $n$ increases, so that we may assume $n$ is small, and therefore $\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\gg 1$. In this limit,
$\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}\int_{0}^{\infty}\frac{x\textrm{d}x}{e^{x}-1}e^{-\frac{x^{2}}{4}\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{3}}}\approx\sqrt{\pi}\left(\frac{E_{c}^{2}}{8n^{4}\pi^{4}\sigma_{\textrm{EM}}^{2}}\right)^{\frac{1}{6}}-1,$ (130)
and so we get
$\Gamma_{\textrm{Schw}}$ $\approx\frac{e^{2}E_{c}^{2}V}{8\sqrt{3}\pi}\sum_{n=1}^{\infty}e^{-3\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{\frac{-1}{3}}}\left(\frac{8\sigma_{\textrm{EM}}^{2}}{n^{2}\pi^{2}E_{c}^{2}}\right)^{\frac{1}{3}}.$ (131)
Substituting the expression for $\sigma_{\textrm{EM}}$ found from eq. 27, and using the definition for $N_{\gamma}^{\textrm{Schw}}$, yields eq. 28. Numerical calculations show that eq. 28 approximates eq. 62 well for $N_{\gamma}\lesssim.1N_{\gamma}^{\textrm{Schw}}$, and continues to be accurate to within an order of magnitude for all $N_{\gamma}<N_{\gamma}^{\textrm{Schw}}$.
Appendix B Behavior of functions of $\chi$